OpenStudy (anonymous):

What is 40% of 3,000? 240 2,400 1,200 120

OpenStudy (anonymous):

1.200

OpenStudy (mokeira):

\[\frac{ 40 }{100 } \times 3000= answer :)\]

OpenStudy (mokeira):

the zeros cancel out

OpenStudy (k_lynn):

Well if 40% equals .4 and of means multiply then the answer should be 1200

OpenStudy (anonymous):

thank u all!!!!

OpenStudy (anonymous):

Your welcome

OpenStudy (mathstudent55):

"percent" means "divide by 100" "of" means "times" (multiplication) "is" means "=" "what" (an unknown) means "x" What is 40% of 3,000 x = 40/100 * 3000 x = 40/100 * 3000 x = 0.4 * 3000 x = 1200